OpenStudy (anonymous):

find the sum to the required number of terms for the geometric series 243+81+27+... to 7 terms i know the formula is s=b1*(1-q^n)/(1-q) but every time i plug it in i get 1093/3 can someone else tell me how to plug it inn please

OpenStudy (anonymous):

b1=243

OpenStudy (anonymous):

q=1/3

OpenStudy (anonymous):

so find it

OpenStudy (anonymous):

n=7

Parth (parthkohli):

\[243 \times \dfrac{1 - \dfrac{1}{3^7}}{1 - \dfrac{1}{3}}\]

OpenStudy (anonymous):

ParthKohli just gave you the explicit answer, where did you mistake

OpenStudy (anonymous):

ok i did that and i still got 1093/3

Parth (parthkohli):

That's weird.

Parth (parthkohli):

You can just calculate \(243 + 81 + 27 + 9 + 3 + 1 + \dfrac{1}{3}\)

OpenStudy (anonymous):

that's correct by the way) xDD

OpenStudy (anonymous):

what did you get when u simplified the equation

Parth (parthkohli):

lol

Parth (parthkohli):

yeah, that is correct.

OpenStudy (anonymous):

1093/3 is correct?

OpenStudy (anonymous):

yes, that's correct) xD

OpenStudy (anonymous):

oh ok thank youu